# Nearly a Decade of Groundbreaking Speckle Interferometry at the International Gemini Observatory

### v1

@@ 4.1.3 Angular Diameters
Because speckle imaging allows the diffraction limits of the telescopes to be achieved, stars with large angular diameters are resolved. ‘Alopeke on Gemini North was used to observe Betelgeuse during its “Great Dimming Event” in February 2020. At the time, Betelgeuse had a $V$ magnitude of 1.4. Figure 5 shows a 562 nm speckle image of the star (Howell et al., 2024, in prep.). Its disk is resolved with an apparent angular diameter of 40 mas.
Figure 5: A 562 nm ‘Alopeke observation of Betelgeuse obtained in February 2020, during the “Great Dimming Event.” Betelgeuse, $V=1.4$, is resolved with an angular diameter of 37 mas. This observation took a total of 12 minutes of Gemini time.
@@ 4.1.4 Stellar Eclipses
Figure 6 shows an 832 nm eclipse observation of ZFT J0220, which allowed the very short, 90-second duration eclipse to be examined in detail. A 562 nm simultaneous light curve was also obtained, providing robust information on the radii of the two stars.
Figure 6: Photometric eclipse light curve of the white dwarf ZTFJ0220 + 2141. The time series consisted of 340$\times$3 second exposures that were used to provide ingress and egress details for the short, 90 second eclipse. Adapted from Scott et al. (2021).
@@ 4.1.5 Evolved Stars and Stellar Remnants
Speckle observations allow for the measurement of winds from Wolf-Rayet stars (Shara et al., 2023), accretion disks in interacting binaries (Figure 7), stellar mergers (Mobeen et al., 2024), nova shells (Figure 8), supernovae (Van Dyk et al., 2024) and well-known specific targets such as R Aqr (Liimets et al. 2025, in prep.) and Eta Carinae (Figure 9).
Figure 7: Speckle image of a hot accretion disk in a nearby interacting binary star. The disk was observed in H-alpha and required 45 minutes of Gemini time.
@@
Figure 9: Two images of the famous star Eta Carina. Left: A false color composite composed of speckle images in the SDSS $u$, 562 nm, and 832 nm filters. These observations used 40 minutes of Gemini time. This image is 0.34 arcseconds on a side. Right: A similar view of the innermost part of the Eta Carinae system, but in the infrared. This Very Large Telescope Interferometer (Weigelt et al., 2016) image was rotated to the same orientation on the sky and zoomed to a similar angular size. The central star and nebular blobs are visible in both images.
@@ 4.2 Solar System Bodies
4.2.1 Angular Diameters
Many objects in our Solar System are large enough to be resolved. Speckle images of Pluto and Charon provided measurements of their diameters, and were the highest resolution images of the (dwarf) planets until the New Horizons fly-by (Howell et al., 2012). Asteroids have also been targets of the Gemini speckle imagers. During its close approach, the asteroid Phaethon (1983TB) was found to be 59 mas in angular size, corresponding to a physical size of 4.1 km (Wooden et al., 2018). Pereira et al. (2023) observed a stellar occultation by the Trans-Neptunian object Quaoar with the aim of improving its shape model and physical parameters and searching for additional material around the body. Figure 10 shows five frames from a 5.3-hour speckle imaging observation sequence of the asteroid 433 Eros. Eros is the second largest near-Earth asteroid. It has an elongated shape and a volume equivalent diameter of $\sim$17 km. This time series covered a single rotation period of the asteroid, and the apparent shape of the non-spherical body is seen changing with time.
@@ 4.2.2 Occultations
Pluto, asteroids, Kuiper Belt Objects, and other Solar System bodies have also had occultations observed by the Gemini speckle imagers, providing simultaneous high-speed two-color photometry (e.g., Sickafoose et al., 2023). Other studies have observed target stars in advance of an occultation to assess whether they are single or multiple (e.g., Schindler et al., 2019).
@@ 4.3.1 Crowded Fields
Figure 11: V4641 Sgr, a $V\sim 14$, high-mass black hole binary, is easily isolated from its very crowded field in our speckle images. Left: A DSS image showing a $4\times 4$ arcminute region around V4641 Sgr (from SIMBAD). Right: Simultaneous SDSS $g$ and SDSS $z$ speckle images for which photometry can be performed without issues of crowding. These images were obtained using ‘Alopeke and 12 minutes of Gemini time.
@@ 4.3.2 Dual Quasars
Observation of a 19th magnitude dual quasar candidate was resulted in the detection of both binary black hole nuclei, and required only 50 minutes of Gemini time (Howell et al., 2021e).
@@ 4.3.3 X-ray Binary Outbursts
Outbursts of X-ray binaries offer opportunities to measure quasi-periodic oscillations during the accretion process, yielding insights into the underlying disks. Tetarenko et al. (2022) used speckle imaging to perform an optical fast timing study of various X-ray binaries. The simultaneous two-color data collected enabled the examination of disk energetics and jet physics. Furthermore, Scott et al. (2018) showed that the ability to detect pulse shape and period in two optical colors at the same time allows precise identification of the pulsational modes involved in the pulsating white dwarf HL Tau.
@@ 5 Summary
The ‘Alopeke and Zorro speckle cameras are permanent visitor instruments at the Gemini North and South telescopes in Hawai’i and Chile. One can request to use these instruments for general queue proposals, DDT and FT proposals, Long and Large Proposals, and soon Target of Opportunity proposals as well. Note that the oversubscription rate of the Gemini telescopes hovers near 3, as compared to 6-7 for HST, Chandra, and JWST.
Optical speckle imaging has the advantage of darker skies (than the IR), no need for AO natural or laser guide stars, inexpensive and simple instrumentation, and easy setup and use. Speckle observations can be made from 400 nm to 1000 nm using, yielding a spectral energy distribution across the optical bandpass for any detected companion. Additionally, speckle imaging is indifferent to whether the target star is single or multiple; the observations and data reduction processes are uniformly applied regardless of the target’s multiplicity. This is not an ability shared by coronagraphic instruments.
@@
As the world of ground-based observational astronomy moves into the era of 30-meter-class telescopes, speckle imagers should be one of the first facility instruments available. Speckle imagers are inexpensive, simple to construct, small in format, and easy to operate, and they obtain fast time series observations and use straightforward software for data reduction. They can also serve as guiders or wavefront sensors. Reaching the optical diffraction limit on a 30-meter-class telescope would provide, at 400 nm, an spectacular angular resolution of 4 mas. This would be an unprecedented, allowing a multitude of new astrophysical science to be performed.

### v2

@@ 6.1.3 Angular Diameters
Because speckle imaging allows the diffraction limit of the telescopes to be achieved, stars with large angular diameters can be resolved. ‘Alopeke on Gemini North was used to observe Betelgeuse during its “Great Dimming Event” in February 2020. At the time, Betelgeuse had a $V$ magnitude of 1.4. Figure 6 shows a 562 nm speckle image of the star along with an unresolved PSF reference star observed near in time (Howell et al., 2025, in prep.). The disk of Betelgeuse is resolved with an apparent angular diameter of 40 mas.
Figure 6: A 562 nm ‘Alopeke observation of Betelgeuse obtained in February 2020, during the “Great Dimming Event.” Betelgeuse, $V=1.4$, is resolved with an angular diameter of 40 mas. The inset shows an unresolved diffraction limited PSF reference star observed near in time to Betelgeuse. This observation took a total of 12 minutes of Gemini time.
@@ 6.1.4 Stellar Eclipses
Figure 7 shows a SDSS i light curve of the r=18 eclipsing double degenerate binary star SDSS J0822+3048 with a period of 40.5 minutes (Kosakowski et al., 2021). Time series standard CCD observations, comprised of 15-second exposures, were simultaneously obtained in SDSS r and SDSS i for about 2 hours, allowing the very short, 90-second duration eclipse to be examined in detail.
Figure 7: Photometric eclipse light curve in SDSS i of the double degenerate binary star SDSS J0822+3048. The time series consisted of 2 hours of 15 second exposures that were used to provide ingress and egress details for the short eclipse. Adapted from Scott et al. (2021).
@@ 6.1.5 Evolved Stars and Stellar Remnants
Speckle observations have allowed the measurement of winds from Wolf-Rayet stars (Shara et al., 2023), shapes of stellar merger remnants (Mobeen et al., 2024), nova shells (Figure 8), supernovae (Van Dyk et al., 2024) and well-known specific targets such as R Aqr (Liimets et al. 2025, in prep.) and Eta Carinae (Figure 9). Eta Car was first observed with speckle imaging by Weigelt and Ebersberger (1986) who found four starlike components in their field of view of 0.8”. The Gemini speckle image shown in (Figure 9) reveals the central light concentration and fans of extended wind emission coming from the binary.
@@
Figure 9: Two images of the famous star Eta Carina. Left: A false color composite composed of speckle images in the SDSS $u$, 562 nm, and 832 nm filters. These observations used a total of 40 minutes of Gemini time. The speckle image is $\sim$0.3 arcseconds (300 mas) on a side. Right: The highest resolution image of Eta Car available today covering the inner 50 mas around the binary. This image was made using the ESO VLTI AMBER instrument in the K band (Weigelt et al., 2016). The orbit of the secondary star around the primary is shown for comparison. The VLTI image contours outline the K band continuum wind emission which is seen to extend to larger distances in the optical speckle composite image.
@@ 6.2 Solar System Bodies
6.2.1 Angular Sizes
Many objects in our Solar System are large enough to be resolved. Speckle images of Pluto and Charon provided measurements of their diameters, and were the highest resolution images of the (dwarf) planets until the New Horizons fly-by (Howell et al., 2012). Asteroids have also been targets of the Gemini speckle imagers. During its close approach, the asteroid Phaethon (1983TB) was found to be 59 mas in angular size, corresponding to a physical size of 4.1 km (Wooden et al., 2018). Pereira et al. (2023) observed a stellar occultation by the Trans-Neptunian object Quaoar with the aim of improving its shape model and physical parameters and searching for additional material around the body. Figure 10 shows five frames from a 5.3-hour speckle imaging observation sequence of the asteroid 433 Eros. Eros is the second largest near-Earth asteroid. It has an elongated shape and a volume equivalent diameter of $\sim$17 km. The time series covered a single rotation period (P=5.3 hr) of the asteroid, and the apparent shape of the non-spherical body is seen to change with time. These data are unpublished and available in the Gemini archive.
@@ 6.2.2 Occultations
Pluto, asteroids, Kuiper Belt Objects, and other Solar System bodies have also had occultations observed by the Gemini speckle imagers, using simultaneous high-speed two-color photometry (e.g., Sickafoose et al., 2023). Other studies have observed target stars in advance of an occultation to assess whether they are single or multiple (e.g., Schindler et al., 2019).
@@ 6.3.1 Crowded Fields
Figure 11: V4641 Sgr, a $V\sim 14$, high-mass black hole binary, is easily isolated from its very crowded field in speckle images. Left: A DSS image showing a $4\times 4$ arcminute region around V4641 Sgr (from SIMBAD). Right: Simultaneous SDSS $g$ and SDSS $z$ speckle images for which photometry can be performed without issues of crowding. These images were obtained using ‘Alopeke and 12 minutes of Gemini time.
@@ 6.3.2 Dual Quasars
Observation of a 19th magnitude dual quasar candidate resulted in the detection of both binary black hole nuclei, and required only 50 minutes of Gemini time (Howell et al., 2021c). Earlier speckle interferometry imaging observations for quasars, even with 6-m telescopes, were limited to brighter ($\leq$16) targets. New speckle observations with Gemini allow nearly diffraction-limited optical imaging opening a new venue to the confirmation and detailed studies of high-z dual quasars.
@@ 6.3.3 X-ray Binary Outbursts
Outbursts of X-ray binaries offer opportunities to measure quasi-periodic oscillations during the accretion process, yielding insights into the underlying disks. Tetarenko et al. (2022) used speckle imaging to perform an optical fast timing study of various X-ray binaries. The simultaneous two-color data collected enabled the examination of disk energetics and jet physics. Furthermore, Scott et al. (2018) showed, using engineering data from the WIYN telescope, that the ability to detect both pulse shape and pulse period in two optical colors at the same time allows precise identification of the pulsational modes involved in the pulsating white dwarf HL Tau.
@@ 7 Summary
The ‘Alopeke and Zorro speckle cameras are permanent visitor instruments at the Gemini North and South telescopes in Hawai’i and Chile. One can request to use these instruments for general queue proposals, DDT and FT proposals, Long and Large Proposals, and Target of Opportunity proposals as well. Note that the oversubscription rate of the Gemini telescopes hovers near 3, as compared to 6-7 for HST, Chandra, and JWST.
Optical speckle imaging has the advantage of darker skies (than the IR), no need for AO natural or laser guide stars, inexpensive and simple instrumentation, and easy setup and use. Gemini speckle observations can be made from 350 nm to 1000 nm, providing the ability to gather a spectral energy distribution across the optical bandpass for any detected companion. Additionally, speckle imaging reaches the diffraction limit and is indifferent to whether the target star is single or multiple; the observations and data reduction processes are uniformly applied regardless of the target’s multiplicity. This is not an ability shared by coronagraphic instruments.
@@
As the world of ground-based observational astronomy moves into the era of 30-40 meter-class telescopes, speckle imagers should be one of the first facility instruments available. Speckle imagers are inexpensive, simple to construct, small in format, light in weight, and easy to operate. The use of EMCCDs and their high-speed imaging ability not only allows speckle imaging to be accomplished but fast time-series photometric observations as well. Such time-series observations of bright stars can also find telescope engineering applications in the analysis of x,y centroids over time to examine guider and tracking errors, and unwanted telescope vibration modes. This type of engineering study has already been successfully implimented at Gemini using ’Alopeke and Zorro. The high-speed imaging cameras can also be used as telescope guiders or wavefront sensors, such as has been done at the WIYN telescope for use with the NEID spectrograph (see e.g., Gupta et al., 2021).
Speckle instrument imagers can also function as standard CCD imagers. Providing, as we have in ’Alopeke and Zorro, an optional wider field of view, the EMCCD cameras can be used as conventional CCD imagers, providing long-exposure photometric images. As a “first-light” instrument, obtaining images throughout the larger final field of view would yield optical quality checks of the delivered PSF across the focal plane.
As speckle interferometry moves to these larger telescopes, reaching the optical diffraction limit on a 30-meter-class telescope would provide, at 400 nm, an spectacular ground-based angular resolution of 4 mas. Such resolution would be unprecedented, allowing a multitude of new astrophysical science cases to be explored.
Author Contributions
All authors listed have made a substantial, direct and intellectual contribution to the work, and approved it for publication. SH & CM-V: Writing – original draft, review and editing. EF, NS, RM, CL, CC, KL, ZH, DC, & SD: review and editing.
Conflict Of Interest
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Generative AI statement
The authors declare that no Generative AI was used in the creation of this manuscript.